OpenStudy (anonymous):

Which represents x2-12x + 36 in factored form?

OpenStudy (smokeysthename):

what # * what # = 36?

OpenStudy (smokeysthename):

and what # + what # = -12 with that being said (x-6)(x-6) is x^2 -12x+36 because x*x=x^2 -6x - 6x = -12x -6*-6 = 36 understand?

OpenStudy (anonymous):

so its (x+6)^2?

OpenStudy (smokeysthename):

no (x-6)^2 (x-6)(x-6) is the same thing as (x-6)^2

OpenStudy (anonymous):

okay thank you

OpenStudy (smokeysthename):

you're welcome - please give me a medal

OpenStudy (anonymous):

Which is the completely factored form of x^6-64y^3?

OpenStudy (smokeysthename):

post a new question